# Scenario-Based Curriculum Generation for Multi-Agent Autonomous Driving

### v1

@@ V Experiments
We conduct two experiments in which we show that MATS-Gym can be used to (1) train multiple agents in an implicit coordination task and (2) for auto-curriculum generation. Both experiments share the base scenario in which agents are tasked with successfully navigating along a route through an intersection in urban environments, similar to the one introduced in the illustrative example. The reward function for this task is a combination of a progress-based and a cruise speed reward:
@@
where $r_{\text{progress}}=p_{t}-p_{t-1}$ is the progress made along the route since the last time-step and $r_{\text{cruise}}=min(\frac{v_{t}}{v_{\text{target}}},1)$ rewards the agent to drive at a cruise velocity $v_{target}$.
For the optimization of the agent policy we use Proximal Policy Optimization (PPO) [25]. The agent receives birdview observations as input and can therefore observe the drivable road area, lane markings, route and other traffic participants.
@@ V-A Learning in Different Action Spaces
In this experiment, we demonstrate our framework’s usability for multi-agent training and discuss how the choice of actions affects the training of cooperative agents.
Scenario Description. Based on the initial scenario, we let four vehicles navigate through a four-way intersection, each with a designated route and a time limit to reach their destination. The starting conditions and goals vary per episode as outlined in Scenic. To succeed, the agent must manage its vehicle, collaborate with other traffic participants, and follow the prescribed route. Termination occurs when all agents are stopped.
@@
Results. Figure 4 shows the learning curves with respect to average episodic reward, collisions and route completions, evaluated over the batch of training data for each policy update. We aggregate the metrics over $k$-size bins of policy updates ($k=5$), reporting the mean and standard deviation.
@@
This experiment emphasizes how the action space design profoundly shapes emergent behavior and affects the learning task’s difficulty. Careful modelling of the problem, encompassing observations and actions, is crucial for multi-agent learning and serves as a key feature in training frameworks. Importantly, it demonstrates that there isn’t a universally optimal level of abstraction for the action space; instead, the choice depends on the problem at hand.
@@ V-B Scenario Based Environment Design
In this experiment, we assess our DCD approach’s ability to align the training scenario distribution with the agent’s capabilities. We examine its impact on agent performance and the distribution of generated levels. Additionally, we compare its effectiveness with PLR and basic domain randomization. We analyze the evolution of sampled scenario parameters for each approach and evaluate their adaptation to the agent’s capabilities, particularly in the early stages of training when the agent struggles with complex scenarios.
@@
  • Assigned route.
@@
  • Target speed for the other vehicles’ controllers.
  • Whether the other vehicles ignores others recklessly.
@@ Results.
Figure 6: Evolution of the parameters that are sampled during the training with DCD (top) and domain randomization (bottom). For each parameter, we report mean and standard deviation over the batch of training data, indicate the direction of major difficulty ($\uparrow$), and the linear interpolation to highlight the general trend (red). We observe that DCD steers the parameter distribution towards configurations that are supposedly easier to solve: straight crossings, fewer and safer NPCs, etc. This leads us to the conclusion that DCD adapts the training distribution faster to the performance level of the agent.
In Figure 5, we compare learning curves over 150K environment steps. Both PLR and DR show similar performance in terms of episodic returns and route completion during training, indicating comparable scenario difficulty. DCD demonstrates a notable increase in training signal, suggesting adaptation by generating progressively easier scenarios in the initial stages of training. We evaluate all approaches on the same 12 hold-out levels, representing different maneuvers with varying numbers of NPCs. Although the performance on the hold-out set is comparable in early stages of the training, DCD and PLR do not suffer as much from performance drops in later stages.
Figure 6 provides insight into the progression of scenario parameters throughout training. DCD shifts the scenario distribution towards less challenging levels, characterized by straight maneuvers for the agent and a low number of vehicles in intersections. For comparison, we also show how DR naturally maintains a uniform distribution over environment parameters.
We also investigate the distribution of level parameters in the replay buffers of PLR and REPAIRED. Figure 7 depicts the average regret associated with various parameter combinations at four distinct checkpoints during the training process. For visualization purposes, we focus on combinations involving the maneuver type and the number of other vehicles (NPCs) present in the intersection. Observing the data, it becomes evident that the average regrets of Dual-Curriculum Design (DCD) levels tend to be higher and exhibit a narrower distribution across fewer parameter combinations compared to those of PLR alone. This trend can be attributed to the presence of the adaptive level sampler, which facilitates expedited convergence towards parameter combinations that are more pertinent to the task at hand.
Our experiments showcase the effectiveness of automatic curriculum design in aligning scenario generation with the agent’s capabilities. Additionally, optimizing the scenario generation process, rather than just the resampling procedure, accelerates the adaptation of the level distribution. Through this experiment, we aim to demonstrate that integrating UED approaches into MATS-Gym facilitates future research in curriculum generation for AD applications.
@@
Figure 7: For each pair of maneuver type and number of NPCs, we average the regrets of all levels with the same parameters in the buffer at four checkpoints throughout the training. We observe that the environment generation policy of DCD leads to more narrowly concentrated regret on fewer level configurations, suggesting faster adaption of the level sampling.
@@ VI Conclusion
In this work, we present MATS-Gym, a multi-agent training framework capable of generating scenario-based auto-curricula for AD tasks in CARLA. MATS-Gym is open to multiple, commonly used frameworks for traffic scenario specification and sampling, including ScenarioRunner, OpenSCENARIO and Scenic. Leveraging Scenic enables sampling from scenario distributions and the integration with UED approaches. This compatibility resembles a promising way to generate more relevant and realistic training scenarios for AD. We demonstrate the usability of our framework in two experiments of multi-agent training and automatic curriculum generation, respectively. By introducing MATS-Gym and demonstrating its application in various experiments, we contribute to ongoing efforts in advancing multi-agent training for autonomous driving.
For future work, we intend to explore the potential of MATS-Gym in adversarial training within multi-agent systems and the integration of generative models for traffic scenario generation.

### v2

@@ V Experiments
We conduct two experiments in which we show that MATS-Gym can be used to: (1) Train multiple agents in an implicit coordination task, and (2) For auto-curriculum generation. Both experiments share the base scenario in which agents are tasked with successfully navigating along a route through an intersection in urban environments, similar to the one introduced in the illustrative example.
The reward function for this task is a combination of a progress-based and a cruise speed reward:
@@
where $r_{\text{progress}}=p_{t}-p_{t-1}$ is the progress made along the route since the last time-step and $r_{\text{cruise}}=min(\frac{v_{t}}{v_{\text{target}}},1)$ rewards the agent to drive at a cruise velocity $v_{target}$. For the optimization of the agent policy we use Proximal Policy Optimization (PPO) [24]. The agent receives birdview observations as input and can therefore observe the drivable road area, lane markings, route and other traffic participants.
@@ V-A Learning in Different Action Spaces
In this experiment, we demonstrate the framework’s usability for multi-agent training and discuss how the choice of actions affects the training of cooperative agents.
Scenario Description. Based on the initial scenario, we let four vehicles navigate through a four-way intersection, each with a designated route and a time limit to reach their destination. The starting conditions and goals vary per episode and are sampled from the description. To succeed, the agent must manage its vehicle, collaborate with other traffic participants, and follow the prescribed route. Termination occurs when all agents are stopped.
@@
Results. Figure 2 shows the learning curves with respect to average episodic reward, collisions and route completions, evaluated over the batch of training data for each policy update. We aggregate the metrics over $k$-sized bins of policy updates ($k=5$), reporting the mean and standard deviation.
@@
This experiment emphasizes how the action space design profoundly shapes emergent behavior and affects the learning task’s difficulty. Careful modelling of the problem, encompassing observations and actions, is crucial for multi-agent learning and serves as a key feature in training frameworks. Importantly, it demonstrates that there is no universally optimal level of abstraction for the action space; instead, the choice depends on the problem at hand.
@@ V-B Scenario Based Environment Design
In this experiment, we assess the ability of the curriculum design algorithm to align the training scenario distribution with the agent’s capabilities. We examine its impact on agent performance and the distribution of generated scenarios. Additionally, we compare its effectiveness with PLR and basic domain randomization. We analyze the evolution of sampled scenario parameters for each approach and evaluate their adaptation to the agent’s capabilities, particularly in the early stages of training when the agent struggles with complex scenarios.
@@
  • Assigned route (straight, left or right turn).
@@
  • Target speed for the other agents.
  • Whether the other vehicles keep safety distances.
@@ Results.
Figure 4: Evolution of the parameter distribution during the training with DCD (top) and domain randomization (bottom). The plot shows the fractions of route types, mean number of NPCs, the fraction of agents that do not keep safety distances and ignore traffic lights and the mean target speed. For each parameter, we report mean and standard deviation over the batch of training data and indicate the direction of major difficulty ($\uparrow$). We observe that DCD steers the parameter distribution towards configurations that are supposedly easier to solve: straight crossings, fewer and safer NPCs, etc. This leads us to the conclusion that DCD adapts the training distribution faster to the performance level of the agent.
@@
Figure 5: For each pair of maneuver type and number of NPCs, we average the regrets of all scenarios with the same parameters in the buffer at four checkpoints throughout the training. We observe that the environment generation policy of DCD leads to more narrowly concentrated regret on fewer configurations, suggesting faster adaption of the scenario sampling.
In Figure 3, we compare learning curves over 150K environment steps. Both PLR and DR show similar performance in terms of episodic returns and route completion during training, indicating comparable scenario difficulty. DCD demonstrates a notable increase in training signal, suggesting adaptation by generating progressively easier scenarios in the initial stages of training. We evaluate all approaches on the same 12 hold-out scenarios, representing different maneuvers with varying numbers of NPCs. Although the performance on the hold-out set is comparable in early stages of the training, DCD and PLR do not suffer as much from performance drops in later stages.
Figure 4 provides insight into the progression of scenario parameters throughout training. DCD shifts the scenario distribution towards less challenging scenarios, characterized by straight maneuvers for the agent and a low number of vehicles in intersections. For comparison, we also show how DR naturally maintains a uniform distribution over environment parameters.
We also investigate the distribution of scenario parameters in the replay buffers of PLR and REPAIRED. Figure 5 depicts the average regret associated with various parameter combinations at four distinct checkpoints during the training process. For visualization purposes, we focus on combinations involving the maneuver type and the number of other vehicles (NPCs) present in the intersection. Observing the data, it becomes evident that the average regrets of Dual-Curriculum Design (DCD) scenarios tend to be higher and exhibit a narrower distribution across fewer parameter combinations compared to those of PLR alone. This trend can be attributed to the presence of the adaptive scenario sampler, which facilitates expedited convergence towards parameter combinations that are more pertinent to the task at hand. Our experiments demonstrate the effectiveness of automatic curriculum design in aligning scenario generation with the agent’s capabilities. Additionally, optimizing the scenario generation process, rather than just the resampling procedure, accelerates the adaptation of the scenario distribution.
@@ VI Conclusion
In this work, we present MATS-Gym, a multi-agent training framework capable of generating scenario-based auto-curricula for AD tasks in CARLA. Leveraging Scenic allows us to sample from scenario distributions and enables the integration with UED approaches. This compatibility resembles a promising way to generate more relevant and realistic training scenarios for AD. We demonstrate the usability of our framework in two experiments of multi-agent training and automatic curriculum generation, respectively. By introducing MATS-Gym and demonstrating its application in various experiments, we contribute to ongoing efforts in advancing multi-agent training for autonomous driving.
VII acknowledgements
This project has received funding from the Austrian FFG ICT of the Future program under grant agreement No 880811.